# Wasserstein Flow Matching: Generative modeling over families of distributions

### v1

@@ Abstract
Generative modeling typically concerns the transport of a single source distribution to a single target distribution by learning (i.e., regressing onto) simple probability flows. However, in modern data-driven fields such as computer graphics and single-cell genomics, samples (say, point-clouds) from datasets can themselves be viewed as distributions (as, say, discrete measures). In these settings, the standard generative modeling paradigm of flow matching would ignore the relevant geometry of the samples. To remedy this, we propose Wasserstein flow matching (WFM), which appropriately lifts flow matching onto families of distributions by appealing to the Riemannian nature of the Wasserstein geometry. Our algorithm leverages theoretical and computational advances in (entropic) optimal transport, as well as the attention mechanism in our neural network architecture. We present two novel algorithmic contributions. First, we demonstrate how to perform generative modeling over Gaussian distributions, where we generate representations of granular cell states from single-cell genomics data. Secondly, we show that WFM can learn flows between high-dimensional and variable sized point-clouds and synthesize cellular microenvironments from spatial transcriptomics datasets. Code is available at WassersteinFlowMatching.
00footnotetext: Equal contribution. Correspondance to doron.haviv12@gmail.com and ap6599@nyu.edu.00footnotetext: Meta was involved only in an advisory role. All experimentation and data processing was conducted at MSKCC.
@@ 1 Introduction
Today’s abundance of data and scalability of training massive neural networks has made it possible to generate hyper-realistic images on the basis of training examples (OpenAI,, 2022), as well as video and audio clips (Vyas et al.,, 2023; Xing et al.,, 2023), and, of course, text (Bubeck et al.,, 2023). All of these are instances of generative modeling: given access to finitely many samples from a distribution, devise a scheme which generates new samples from the same distribution. Generative modeling has also been revolutionary in the biomedical sciences, for drug design (Jumper et al.,, 2021) and single-cell genomics (Lopez et al.,, 2018). Nearly all frameworks exploit the notion that datasets (of, say, genomic profiles of cells, images, videos, or corpora of text documents) are instantiations of probability measures, and the task is to transform a point sampled from random noise to generate a data point that obeys the distribution of interest.
Among the zoo of available generative models, one approach noted for its flexibility and simplicity is Flow Matching (FM) (Albergo and Vanden-Eijnden,, 2022; Lipman et al.,, 2022; Liu et al.,, 2022). For a fixed target probability measure, FM learns an implicitly defined vector field that can transform a source measure (e.g., the standard Gaussian) to the target. Unlike discrete time and probabilistic generative models (such as Denoising Diffusion Models by Song et al., (2020)), FM learns a deterministic, continuous normalizing flow by regressing onto a simple conditional probability flow. This approach, while originally designed for Euclidean domains, can be readily adopted to Riemannian geometries (Chen and Lipman,, 2023). Riemannian flow matching is widely used for generating samples over geometries such as spheres, tori, translation/rotation groups, simplices, triangular meshes, mazes, and molecular positions and structures.
The Wasserstein geometry, a canonical geometry over distributions, does not easily fit into any of these existing frameworks and has not been successfully adapted for flow matching. This geometry is useful, for example, in computational graphics where collections of 3D shapes are represented as empirical distributions (point-clouds). Likewise, recent developments in single-cell genomics analysis have demonstrated that gene-expression profiles from groups of cells aggregated via their mean and covariance can capture cellular microenvironments or highlight fine-grain clusters (Haviv et al., 2024b, ; Persad et al.,, 2023). For both point-cloud and Gaussian settings, it is natural to search for a unified generative model that respects the underlying geometry of the data, namely, treating each sample as itself a probability distribution.
@@
We introduce Wasserstein Flow Matching (WFM), a principled extension of the FM framework lifted to the space of probability distributions. As illustrated in Figure 1, a single point in our source and target datasets is itself a distribution (e.g., a single discrete measure or a single Gaussian), and our aim is to learn vector fields acting on the space of probability distributions and match the optimal transport map, which is the geodesic in Wasserstein space. WFM is an instantiation of Riemannian FM  (Chen and Lipman,, 2023), where we train a neural model to learn a continuous normalizing flow (CNF) between distributions over distributions.
@@
[…]
Figure 1: Left: Table contrasting FM methods over $\mathbb{R}^{d}$, general manifolds $\mathcal{M}$, categorical and Dirichlet distributions on the $d$-simplex $\Delta_{d}$, and finally, our approach, FM problems defined over $\mathcal{P}(\mathbb{R}^{d})$. Right: WFM overview, which learns flows over distributions over distributions.
We demonstrate the effectiveness of our approach for generative modeling between distributions over Gaussian distributions and distributions over point-clouds. The former task is motivated by recent directions in single-cell and spatial transcriptomics (Haviv et al., 2024b, ; Persad et al.,, 2023), where we consider matching problems over the Bures–Wasserstein space (BW), the Gaussian submanifold of the Wasserstein space. In this case, we show that WFM can be further modified, resulting in the Bures–Wasserstein FM (BW-FM) algorithm. We validate BW-FM on a variety of Gaussian-based datasets, where we observe that samples generated by our algorithm are significantly more robust than naïve approaches which do not fully exploit the underlying geometry of the data. In turn, we present a generative model for cell states and niches from single-cell genomics data.
Point-cloud generation is made possible by two distinct, yet crucial, algorithmic primitives: (1) incorporating transformers in our neural network architecture (Vaswani,, 2017; Lee et al.,, 2019), and (2) recent algorithmic advances in entropic optimal transport (Pooladian and Niles-Weed,, 2021). Indeed, our WFM algorithm performs generative modeling in the Wasserstein space, where geodesics are given by pushforwards of optimal transport (OT) maps; see Section 2.3 for more information. Both the transformer architecture and entropic optimal transport are crucial to approximating the OT map between independent point-clouds. Indeed, the permutation equivariance of attention makes the transformer a natural basis for our model, inherently modeling the equivariance feature of the Wasserstein geometry while maintaining scalability in high-dimensions.
For datasets of 3D point-clouds with uniform sizes, the performance of WFM is comparable to other current generative models. However, due to their particular training paradigms (namely the voxelization of 3D spaces), contemporary approaches cannot scale to high-dimensional point-clouds and fail on datasets with variable sized examples. Conversely, WFM succeeds in the high-dimensional and inhomogeneous settings, unlocking generative modeling to new, previously uncharted domains such as synthesizing niches from spatial genomics data. The ability to model tissue biology in this generative manner could enhance our understanding of how environment is associated with cell state. In the context of many diseases, most notably cancer and its tumor-immune microenvironment, these insights are critical for developing novel therapeutics (Binnewies et al.,, 2018).
@@ 2 Background and related work
We let $\mathcal{P}_{2}(\mathbb{R}^{d})$ denote the set of probability distributions over $\mathbb{R}^{d}$ with finite second moment, and write $\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$ to be those with densities. For a probability measure $\mu$ and (vector-valued) function $f$, we interchangeably write $\int\|f(x)\|^{2}\,\mathrm{d}\mu(x)$ and $\|f\|^{2}_{L^{2}(\mu)}$. Let $\mathcal{M}$ be a Riemannian manifold, with $\mathcal{P}(\mathcal{M})$ defining the space of probability measures over said manifold. For $x\in\mathcal{M}$, we write $\mathcal{T}_{x}\mathcal{M}$ to mean the tangent space of the manifold at $x$, and write the metric on the tangent space (at $x$) as $g(x)$. For $x_{0}\in\mathcal{M}$ with initial velocity $v\in\mathcal{T}_{x_{0}}\mathcal{M}$, the terminal location of the resulting geodesic is expressed as the output of the exponential map $v\mapsto\exp_{x_{0}}(v)\in\mathcal{M}$. Similarly, for an initial point $x_{0}$ and terminal location $x_{1}$, the logarithmic map defines the tangent vector, denoted $x_{1}\mapsto\log_{x_{0}}(x_{1})$, such that $\exp_{x_{0}}(\log_{x_{0}}(x_{1}))=x_{1}$. The set of symmetric matrices (resp. positive definite matrices) over $\mathbb{R}^{d}$ are denoted by $\mathbb{S}^{d}$ (resp. $\mathbb{S}^{d}_{++}$).
@@ 2.1 Riemannian flow matching
We first briefly discuss the Riemannian flow matching (RFM) framework of Chen and Lipman, (2023). Let $\mathfrak{p}_{0}$ be the source distribution and $\mathfrak{p}_{1}$ be the target distribution over a Riemannian manifold $\mathcal{M}$, and let $(\gamma_{t})_{t\in[0,1]}$ be a curve of probability measures satisfying $\gamma_{0}=\mathfrak{p}_{0}$ and $\gamma_{1}=\mathfrak{p}_{1}$. Letting $(w_{t})_{t\in[0,1]}$ denote a family of vector fields, we say that the pair $(\gamma_{t},w_{t})_{t\in[0,1]}$ satisfy the continuity equation with respect to the metric $g$, abbreviated to $(\gamma_{t},w_{t})\in\mathfrak{C}_{g}$ if
@@
assuming access to a pair $(\gamma_{t},w_{t})_{t\in[0,1]}$ that satisfies equation 1, which is not possible in many scenarios. Borrowing insights from recent work (e.g., Albergo and Vanden-Eijnden, (2022); Lipman et al., (2022); Liu et al., (2022)), the authors construct a simple vector field that satisfies the continuity equation, resulting in the tractable objective
@@
where, for example, $x_{t}=\exp_{x}((1-t)\log_{x}(y))\in\mathcal{M}$, and $\dot{x}_{t}\in\mathcal{T}_{x_{t}}\mathcal{M}$. For complete discussions and proofs, see (Chen and Lipman,, 2023). Once $f_{\theta}$ is appropriately fit using equation 2, we can generate new samples from $\mathfrak{p}_{1}$: start by sampling $X_{0}\sim\mathfrak{p}_{0}$, then follow $\dot{X}_{t}=f_{\theta}(X_{t},t)$ numerically by discretizing the dynamics given by the exponential map, resulting in $X_{1}\sim\mathfrak{p}_{1}$. We emphasize that the dynamics are only simulated at inference time and not when training $f_{\theta}$, which is commonly known as a simulation-free training paradigm.
@@ 2.2 Related work
Generative models for point-clouds.
Paralleling the progress in generative models for natural images, the field of point-cloud generation is rapidly expanding. Many different models have been used from this task, namely generative-adversarial-nets (Achlioptas et al.,, 2018), variational autoencoders (Gadelha et al.,, 2018), normalizing flows (Yang et al.,, 2019; Kim et al.,, 2020; Klokov et al.,, 2020), diffusion (Zhou et al.,, 2021; Cai et al.,, 2020) and even euclidean FM (Wu et al.,, 2023). Thus far, these approaches are limited to uniformly sized point-clouds in 2D & 3D, and fail on high-dimensional spaces which cannot be voxelized.
@@
Our work is not the first to instantiate Riemannian FM with a manifold of probability measures. Two notable works are Fisher FM (Davis et al.,, 2024) and Categorical FM (Cheng et al.,, 2024), which consider the FM algorithm with respect to the Fisher–Rao geometry Amari, (2016); Nielsen, (2020) over the $d$-dimensional simplex $\Delta_{d}$. The work of Stark et al., (2024) is similar in spirit, where they focus on the Dirichlet distribution for generation of discrete data. Another related work is that of Atanackovic et al., (2024), called Meta FM. Their approach requires pairs of distributions which are already coupled, with the goal of solving FM between a distribution over pairs. In contrast, we emphasize that our proposed Wasserstein FM applies between two separate uncoupled distributions over distributions.
@@ 2.3 Wasserstein geometry
The (squared) $2$-Wasserstein distance between two probability measures $\mu,\nu\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$ is given by the non-convex optimization problem over vector-valued maps $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$
@@
where the pushforward constraint, written $T_{\sharp}\mu=\nu$, means that, for $X\sim\mu$, the image follows $T(X)\sim\nu$. The minimizer to equation 3 is called the optimal transport (OT) map, denoted $T^{\mu\to\nu}_{\star}$ (we abbreviate this to $T_{\star}$ when it is clear from context). The existence and uniqueness of the optimal transport map under the stated regularity conditions is due to Brenier, (1991).
The Wasserstein space is the space of probability densities with finite second moment endowed with the Wasserstein distance; this space is known to be a metric space (Villani,, 2009). Following the celebrated work of Otto, (2001), the Wasserstein space can be formally (meaning, non-rigorously) viewed as a Riemannian manifold, whose properties we now describe in brief; see e.g., Ambrosio et al., (2008) for a rigorous treatment.
Following the definition in Theorem 8.5.1 from Ambrosio et al., (2008), the tangent space at a point $\mu\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$ consists of all possible tangent vectors that emanate from $\mu$, written formally as
$\mathcal{T}_{\mu}\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})\coloneqq\overline{\{\lambda(T_{\star}^{\mu\to\nu}-\mathrm{id})\ :\ \lambda>0,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})\}}^{L^{2}(\mu)}\,,$
where the overline denotes the closure of the set in $L^{2}(\mu)$, and the norm on the tangent space is also $L^{2}(\mu)$. The exponential and logarithmic maps read
$v\mapsto\exp_{\mu}(v)\coloneqq(\mathrm{id}+v)_{\sharp}\mu\,,\quad\nu\mapsto\log_{\mu}(\nu)\coloneqq T_{\star}^{\mu\to\nu}-\mathrm{id}\,,$
@@
$\mu_{t}\coloneqq(T_{t}^{\mu\to\nu})_{\sharp}\mu\coloneqq((1-t)\text{id}+tT_{\star}^{\mu\to\nu})_{\sharp}\mu\equiv\exp_{\mu}((1-t)\log_{\mu}(\nu))\,,$ (4)
where the last expression writes the pushforward in terms of the exponential and logarithmic maps. Equivalently, at the level of the random variables, one can write $X_{t}=(1-t)X_{0}+tT_{\star}^{\mu\to\nu}(X_{0})$, where $X_{0}\sim\mu$ and $X_{t}\sim\mu_{t}$ for any $t\in[0,1]$. Combined with $(v_{t})_{t\in[0,1]}$ a suitable family of vector fields, the McCann interpolation satisfies the continuity equation equation 1 over $\mathbb{R}^{d}$, re-written as
@@
where the divergence operator is the usual Euclidean one over $\mathbb{R}^{d}$, thus we write $(\mu_{t},v_{t})\in\mathfrak{C}$. The link between the constant speed geodesics and the $2$-Wasserstein distance can be viewed from the celebrated Benamou–Brenier formulation of optimal transport (Benamou and Brenier,, 2000):
@@
The vector field equation 7 should be interpreted as the time-derivative of the McCann interpolation:
$\dot{X}_{t}=(T_{\star}^{\mu\to\nu}-\text{id})(X_{0})=(T_{\star}^{\mu\to\nu}-\text{id})\circ(T_{t}^{\mu\to\nu})^{-1}(X_{t})\,,\qquad X_{0}\sim\mu\,.$

### v2

@@ Abstract
Generative modeling typically concerns transporting a single source distribution to a target distribution via simple probability flows. However, in fields like computer graphics and single-cell genomics, samples themselves can be viewed as distributions, where standard flow matching ignores their inherent geometry. We propose Wasserstein flow matching (WFM), which lifts flow matching onto families of distributions using the Wasserstein geometry. Notably, WFM is the first algorithm capable of generating distributions in high dimensions, whether represented analytically (as Gaussians) or empirically (as point-clouds). Our theoretical analysis establishes that Wasserstein geodesics constitute proper conditional flows over the space of distributions, making for a valid FM objective. Our algorithm leverages optimal transport theory and the attention mechanism, demonstrating versatility across computational regimes: exploiting closed-form optimal transport paths for Gaussian families, while using entropic estimates on point-clouds for general distributions. WFM successfully generates both 2D & 3D shapes and high-dimensional cellular microenvironments from spatial transcriptomics data. Code is available at Wasserstein Flow Matching.
Machine Learning, ICML
@@ 1 Introduction
Today’s abundance of data and scalability of training massive neural networks has made it possible to generate hyper-realistic images on the basis of training examples (OpenAI, 2022), as well as video and audio clips (Vyas et al., 2023; Xing et al., 2023), and, of course, text (Bubeck et al., 2023). All of these are instances of generative modeling: given access to finitely many samples from a distribution, devise a scheme which generates new samples from the same distribution. Generative modeling has also been revolutionary in the biomedical sciences, for drug design (Jumper et al., 2021), and single-cell genomics (Lopez et al., 2018). Nearly all frameworks exploit the notion that datasets (of, say, genomic profiles of cells, images, videos, or corpora of text documents) are instantiations of probability measures, and the task is to transform a point sampled from random noise to generate a data point that obeys the distribution of interest.
Among the zoo of available generative models, one approach noted for its flexibility and simplicity is Flow Matching (FM) (Albergo & Vanden-Eijnden, 2022; Lipman et al., 2022; Liu et al., 2022). For a fixed target probability measure, FM learns an implicitly defined vector field that can transform a source measure (e.g., the standard Gaussian) to the target. Unlike discrete time and probabilistic generative models (such as Diffusion Models by Song et al. (2020)), FM learns a deterministic, continuous normalizing flow by regressing onto a simple conditional probability flow. This approach, while originally designed for Euclidean domains, can be readily adopted to Riemannian geometries (Chen & Lipman, 2023). Riemannian flow matching (RFM) is widely used for generating samples over geometries such as spheres, tori, translation/rotation groups, simplices, triangular meshes, mazes, and molecular positions and structures.
However, modern data can exhibit a richer structure in which the samples themselves are distributions. In computational graphics, 3D models are shape distributions observed through point-clouds. Likewise, recent developments in single-cell genomics analysis have demonstrated that gene-expression profiles from groups of cells aggregated via their mean and covariance can capture cellular microenvironments or highlight fine-grain clusters (Haviv et al., 2024b; Persad et al., 2023). For both general distributions and Gaussian settings, it is natural to search for a unified generative model that respects the underlying geometry of the data.
@@
To this end, we introduce Wasserstein Flow Matching (WFM), a principled extension of the FM framework lifted to the space of probability distributions. As illustrated in Figure 1, a single point in our source and target datasets is itself a distribution. These distributions are either represented analytically as Gaussians or realized through point-cloud samples. Our aim is to learn vector fields acting on the space of probability distributions and match the optimal transport map, which is the geodesic in Wasserstein space. WFM is an instantiation of Riemannian FM  (Chen & Lipman, 2023), where we train a neural model to learn a continuous normalizing flow (CNF) between distributions over distributions.
We demonstrate the effectiveness of our approach for generative modeling between distributions over Gaussian distributions and distributions over point-clouds, which are realization of general distributions. The former task is motivated by recent directions in single-cell and spatial transcriptomics (Haviv et al., 2024b; Persad et al., 2023), where we consider matching problems over the Bures–Wasserstein space (BW), the Gaussian submanifold of the Wasserstein space. In this case, we show that WFM can be further modified, resulting in the Bures–Wasserstein FM (BW-FM) algorithm. We validate BW-FM on a variety of Gaussian-based datasets, where we observe that samples generated by our algorithm are significantly more robust than naïve approaches which do not fully exploit the underlying geometry of the data. As an application, we present a generative model for cell states and niches from single-cell genomics data.
@@
Table 1: Contrasting FM methods over $\mathbb{R}^{d}$, general manifolds $\mathcal{M}$, categorical and Dirichlet distributions on the $d$-simplex $\Delta_{d}$, and finally, our approach, FM problems defined over $\mathcal{P}(\mathbb{R}^{d})$.
Generation of general distributions is made possible by two distinct, yet crucial, algorithmic primitives: (1) incorporating transformers in our neural network architecture (Vaswani, 2017; Lee et al., 2019), and (2) recent algorithmic advances in entropic optimal transport (Pooladian & Niles-Weed, 2021). Our WFM algorithm performs generative modeling in the Wasserstein space, where geodesics are given by pushforwards of optimal transport (OT) maps; see Section 2.3 for more information. Since these OT maps lack closed-form solutions for general distributions, we represent distributions as point-clouds and estimate the maps using entropic optimal transport. The permutation equivariance of attention makes transformers a natural basis for our model, inherently capturing the equivariance of Wasserstein geometry while maintaining scalability in high dimensions.
For datasets of distributions in 3D, WFM’s performance matches existing generative models. However, due to their reliance on voxelization, current approaches cannot scale to high-dimensional distributions and fail when distributions are realized as point-clouds with variable sizes. Conversely, WFM succeeds in these challenging settings, enabling generative modeling in new domains like synthesizing tissue microenvironments from spatial genomics data. Modeling tissue biology in this generative manner could enhance our understanding of how environment is associated with cell state. In the context of many diseases, most notably cancer and its immune microenvironment, these insights are critical for developing novel therapeutics (Binnewies et al., 2018).
Refer to caption $\mathfrak{p}_{0}$$\mathfrak{p}_{1}$$\mu$$\nu$$\mathcal{N}(m_{\mu},\Sigma_{\mu})$$\mathcal{N}(m_{\nu},\Sigma_{\nu})$$\tfrac{1}{m}\sum_{i}\delta_{x_{i}}$$\tfrac{1}{n}\sum_{j}\delta_{y_{j}}$
Figure 1: WFM learns flows between distributions over distributions, where distributions are either analytically represented (Gaussians) or empirically observed through point-clouds.
@@ 2 Background and related work
We let $\mathcal{P}_{2}(\mathbb{R}^{d})$ denote probability distributions over $\mathbb{R}^{d}$ with finite second moment, and write $\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$ for those with densities. For a probability measure $\mu$ and function $f$, we write $|f|^{2}_{L^{2}(\mu)}$ for the squared $L^{2}(\mu)$ norm. Let $\mathcal{M}$ be a Riemannian manifold with tangent space $\mathcal{T}_{x}\mathcal{M}$ at $x$. For $x_{0}\in\mathcal{M}$ with velocity $v\in\mathcal{T}_{x_{0}}\mathcal{M}$, the exponential map $\exp_{x_{0}}(v)$ gives the terminal location, while the logarithmic map $\log_{x_{0}}(x_{1})$ gives the initial velocity for the geodesic from $x_{0}$ to $x_{1}$. The set of symmetric (resp. positive definite) matrices over $\mathbb{R}^{d}$ are denoted by $\mathbb{S}^{d}$ (resp. $\mathbb{S}^{d}_{++}$).
@@ 2.1 Riemannian flow matching
We first briefly discuss the Riemannian flow matching (RFM) framework of Chen & Lipman (2023). Let $\mathfrak{p}_{0}$ be the source distribution and $\mathfrak{p}_{1}$ be the target distribution over a Riemannian manifold $\mathcal{M}$, and let $(\gamma_{t})_{t\in[0,1]}$ be a curve of probability measures satisfying $\gamma_{0}=\mathfrak{p}_{0}$ and $\gamma_{1}=\mathfrak{p}_{1}$. Letting $(w_{t})_{t\in[0,1]}$ denote a family of vector fields, we say that the pair $(\gamma_{t},w_{t})_{t\in[0,1]}$ satisfy the continuity equation with respect to the metric $g$, abbreviated to $(\gamma_{t},w_{t})\in\mathfrak{C}_{g}$ if
@@
assuming access to a pair $(\gamma_{t},w_{t})_{t\in[0,1]}$ that satisfies (1). This is not possible in many scenarios. Borrowing insights from recent work (e.g., Albergo & Vanden-Eijnden (2022); Lipman et al. (2022); Liu et al. (2022)), the authors construct a simple vector field that satisfies the continuity equation, resulting in the tractable objective
@@
where, for example, $x_{t}=\exp_{x}((1-t)\log_{x}(y))\in\mathcal{M}$, and $\dot{x}_{t}\in\mathcal{T}_{x_{t}}\mathcal{M}$. For complete discussions and proofs, see Chen & Lipman (2023, Section 3.1). Once $f_{\theta}$ is appropriately fit using (2), we can generate new samples from $\mathfrak{p}_{1}$: start by sampling $X_{0}\sim\mathfrak{p}_{0}$, then follow $\dot{X}_{t}=f_{\theta}(X_{t},t)$ numerically by discretizing the dynamics given by the exponential map, resulting in $X_{1}\sim\mathfrak{p}_{1}$. We emphasize that the dynamics are only simulated at inference time and not when training $f_{\theta}$, commonly known as a simulation-free training paradigm.
@@ 2.2 Related work
Generative models for shapes.
Paralleling the progress in generative models for natural images, the field of shape generation is rapidly expanding. Many different models have been used from this task, namely generative-adversarial-nets (Achlioptas et al., 2018), variational autoencoders (Gadelha et al., 2018), normalizing flows (Yang et al., 2019; Kim et al., 2020; Klokov et al., 2020), diffusion (Zhou et al., 2021; Cai et al., 2020) and even euclidean FM (Wu et al., 2023a). Thus far, these approaches are limited to uniform-sized samplings of shape distributions in 2D & 3D, and fail on datasets where realizations are variably sized.
Figure 2: In the presence of sufficiently many samples, all methods generate Gaussians along the whole spiral, and our Riemannian BW-FM algorithm produces the most consistent samples. Other methods produce Gaussians with degenerate covariance, as they do not model geometry of the data. When there are only few examples, BW-FM accurately reconstructs the training data; see Figure S1.
@@
Our work is not the first to instantiate Riemannian FM with a manifold of probability measures. Two notable works are Fisher FM (Davis et al., 2024) and Categorical FM (Cheng et al., 2024), which consider the FM algorithm with respect to the Fisher–Rao geometry (Amari, 2016; Nielsen, 2020) over the $d$-dimensional simplex $\Delta_{d}$. The work of Stark et al. (2024) is similar in spirit, where they focus on the Dirichlet distribution for generation of discrete data. Another related work is that of Atanackovic et al. (2024), called Meta FM. Their approach requires pairs of distributions which are already coupled, with the goal of solving FM between a distribution over pairs. In contrast, we emphasize that our proposed Wasserstein FM applies between two separate uncoupled distributions over distributions.
Generative models for single-cell genomics.
Deep learning based generative models have transformed single-cell genomics through various approaches. Variational auto-encoders (Lopez et al., 2018; Gayoso et al., 2022) have successfully addressed technical artifacts, integrating multi-modal data and imputing missing features in scRNA-seq data. More recently, Transformer-based foundation models have been noted for their ability to integrate large atlases of data (Cui et al., 2024; Theodoris et al., 2023). Flow Matching has also emerged as a promising direction (Klein et al., 2024; Eyring et al., 2024) for learning both balanced and unbalanced OT maps between cell populations. While these prior FM applications focus on cell-to-cell mappings, our work introduces a new paradigm: generating entire distributions representing cellular populations. This is particularly relevant for spatial genomics, where cellular microenvironments are naturally represented as distributions. WFM enables synthesis of whole cellular neighborhoods, a novel approach in generative modeling for the single-cell field.
@@ 2.3 Wasserstein geometry
The (squared) $2$-Wasserstein distance between two probability measures $\mu,\nu\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$ is given by the optimization problem over vector-valued maps $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$
@@
where the pushforward constraint, written $T_{\sharp}\mu=\nu$, means that for $X\sim\mu$, $T(X)\sim\nu$. The minimizer to (3) is called the optimal transport (OT) map, denoted $T^{\mu\to\nu}_{\star}$ (we abbreviate this to $T_{\star}$ when clear). The existence and uniqueness of the optimal transport map under the stated regularity conditions is due to Brenier (1991).
The Wasserstein space is the space of probability densities with finite second moment endowed with the Wasserstein distance; this space is known to be a metric space (Villani, 2009). Following the celebrated work of (Otto, 2001), the Wasserstein space can be formally (meaning, non-rigorously) viewed as a Riemannian manifold, whose properties we now describe in brief; see e.g., (Ambrosio et al., 2008) for a rigorous treatment. Following the definition by Ambrosio et al. (2008, Theorem 8.5.1), the tangent space at a point $\mu\in\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})$ consists of all possible vectors that emanate from $\mu$, written as
$\mathcal{T}_{\mu}\mathcal{P}_{2,\text{ac}}(\mathbb{R}^{d})\!\coloneqq\!\overline{\{\lambda(T_{\star}^{\mu\to\nu}-\mathrm{id})\!:\!\lambda>0,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})\}}^{L^{2}(\mu)}$
where the overline denotes the closure of the set (i.e., the set and its limit points) in $L^{2}(\mu)$, and the norm on the tangent space is $L^{2}(\mu)$. The exponential and logarithmic maps read
$v\mapsto\exp_{\mu}(v)\coloneqq(\mathrm{id}+v)_{\sharp}\mu\,,\,\nu\mapsto\log_{\mu}(\nu)\coloneqq T_{\star}^{\mu\to\nu}-\mathrm{id}\,,$
@@
$\mu_{t}$ $\coloneqq(T_{t}^{\mu\to\nu})_{\sharp}\mu\coloneqq((1-t)\text{id}+tT_{\star}^{\mu\to\nu})_{\sharp}\mu$ (4)
$\equiv\exp_{\mu}(t\log_{\mu}(\nu))\,,$
where the last expression writes the pushforward in terms of the exponential and logarithmic maps. Equivalently, at the level of the random variables, one can write $X_{t}=(1-t)X_{0}+tT_{\star}^{\mu\to\nu}(X_{0})$, where $X_{0}\sim\mu$ and $X_{t}\sim\mu_{t}$ for any $t\in[0,1]$. Combined with $(v_{t})_{t\in[0,1]}$ a suitable family of vector fields, the McCann interpolation satisfies the continuity equation (1) over $\mathbb{R}^{d}$, re-written as
@@
where the divergence operator is the usual Euclidean one over $\mathbb{R}^{d}$, thus we write $(\mu_{t},v_{t})\in\mathfrak{C}$. The link between the constant speed geodesics and the $2$-Wasserstein distance can be viewed from the celebrated Benamou–Brenier formulation of optimal transport (Benamou & Brenier, 2000):
@@
The vector field (7) should be interpreted as the time-derivative of the McCann interpolation, with $X_{0}\sim\mu$
$\dot{X}_{t}$ $=(T_{\star}^{\mu\to\nu}-\text{id})(X_{0})$
$=(T_{\star}^{\mu\to\nu}-\text{id})\circ(T_{t}^{\mu\to\nu})^{-1}(X_{t})\,.$